People visiting zoo in the morning to the evening are in the ratio 5:3 , if 48 more people visit the zoo in the morning than in the evening than in the evening ,how many people visit zoo in the morning ?
1
Expert's answer
2020-03-18T17:58:27-0400

Let's say x people visit zoo in the evening than

"\frac {48\ +\ x}{x}\ =\ \frac 5 3 \newline
48 + x\ =\ \frac 5 3 x \newline
\frac 2 3 x\ =\ 48 \newline
x\ =\ 48 * \frac 3 2 = 72 \newline
48 + x = 48 + 72 = 120"

Answer: 120 people visit zoo in the morning.